A 10% down payment for a house is RM10,000. The balance has to be repaid by making monthly instalments for ten years. If the interest rate charged is 6% per annum on reducing balance, calculate, using the Constant Ratio formula, (a) the cash price of the house, (b) the total interest charged, (c) the instalment price, and (d) the monthly payment


1
Expert's answer
2021-01-13T19:31:17-0500

(a) the cash price of the house is:

"10,000\/0.1 = RM100,000.10,000\/0.1=RM100,000."

(b) the total interest charged is:

"90,000\u00d70.06\u00d710 = RM54,000.90,000\u00d70.06\u00d710=RM54,000."

(c) the instalment price is:

"10,000+90,000+54,000 = RM154,000.10,000+90,000+54,000=RM154,000."

(d) the monthly payment is:

"P = \frac{154,000} {12×10} = RM1283.33."